### v1

# Enhancing Fault Detection for Large Language Models via Mutation-Based Confidence Smoothing

Abstract.
Large language models (LLMs) achieved great success in multiple application domains and attracted huge attention from different research communities recently. Unfortunately, even for the best LLM, there still exist many faults that LLM cannot correctly predict. Such faults will harm the usability of LLMs. How to quickly reveal them in LLMs is important, but challenging. The reasons are twofold, 1) the heavy labeling effort for preparing the test data, and 2) accessing closed-source LLMs such as GPT4 is money-required. To handle this problem, in the traditional deep learning testing field, test selection methods have been proposed for efficiently testing deep learning models by prioritizing faults. However, the usefulness of these methods on LLMs is unclear and under exploration. In this paper, we first study the effectiveness of existing fault detection methods for LLMs. Experimental results on four different tasks (including both code tasks and natural language processing tasks) and four LLMs (e.g., LLaMA and GPT4) demonstrated that existing fault detection methods cannot perform well on LLMs (e.g., seven out of eight methods perform worse than random selection on LLaMA). To enhance existing fault detection methods, we propose MuCS, a prompt Mutation-based prediction Confidence Smoothing method for LLMs. Concretely, we mutate the prompts and compute the average prediction confidence of all mutants as the input of fault detection methods. The results show that our proposed solution significantly enhances existing methods with the improvement of test relative coverage by up to 97.64%.
1. Introduction
Since the last few years, the potential of large language models (LLMs) has brought hope to general artificial intelligence (AGI). LLMs achieved promising and even the best results compared to other methods in many areas, such as question answering (Zhou et al., 2022), sentiment analysis (Zhang et al., 2023a), and code understanding (Ma et al., 2023) that make LLMs become the first choice of deep learning models in such tasks. Besides the most famous application ChatGPT 111OpenAI, ChatGPT, https://chat.openai.com/, multiple LLMs have been proposed, e.g., LLaMA (Touvron et al., 2023), Alpaca (Taori et al., 2023), and Cerebras-GPT (Dey et al., 2023). More interestingly, focusing on software engineering (SE) tasks, a series of code-related LLMs with good programming ability have been released to help the software development such as StarCoder (Li et al., 2023a) and CodeLLaMA (Roziere et al., 2023). Learning to use LLMs is already a basic skill for us in daily life.
Despite the success, the same as other types of deep learning models, LLMs also suffer from different problems that limit their practical usage. For example, LLMs sometimes generate unrelated outputs to the inputs or incorrect results misaligned with established world knowledge, which is known as the hallucination problem (Zhang et al., 2023b). Besides, recent work (Dong et al., 2023a) showed that LLMs can be easily fooled by adversarial attacks and are not robust. Furthermore, considering SE tasks, existing code-related LLMs cannot capture equivalent semantics when the given natural language prompt expresses the same meaning in different languages (Peng et al., 2024). Those limitations remind us it is necessary to comprehensively evaluate and test the capabilities of LLMs before deployment.
The basic way to test LLMs is to prepare as much as possible test data to assess the performance of LLMs accordingly. However, preparing such test data is heavy work due to the complicated process of data collection, data cleaning, and data labeling. Especially, data labeling requires human effort with domain knowledge which is the most challenging part. Moreover, some LLMs are closed-source which requires funding expenses, e.g., it tasks $0.03/1K tokens of inputs for GPT4 using OpenAI API. Therefore, it is almost impossible to test LLMs for people who have limited budgets. How to efficiently test the performance of LLMs with less effort becomes an urgent problem.
To tackle the data labeling issue, in the field of deep learning testing, multiple fault detection methods that quickly identify fault data (data the model has incorrect prediction) in the test set without using the labeling information (Hu et al., 2024) have been proposed. Based on the information required, existing methods can be divided into learning-based methods and output-based methods. Those methods have been widely studied in the classical deep learning models and have proven to be effective in saving the labeling budget during testing. However, there are no works that explore the usefulness of fault detection methods for LLMs. It is unclear whether we can directly employ such methods to test LLMs and save our effort or not.
To bridge this gap, in this paper, we first study the effectiveness of existing fault detection methods on LLMs. The major challenge in conducting such a study is that most existing fault detection methods are proposed for classification tasks and use the output probabilities for the selection. It is necessary to design the prompt to guide LLMs to produce the output with their confidence. To solve this, for datasets that have more than two classes, we add examples in each class in the prompt as guidance to ask LLMs to predict new data. In total, our study covers nine fault detection methods including both learning-based methods, e.g., TestRank (LI et al., 2021), and output-based methods, e.g., ATS (Gao et al., 2022). Based on the experiment results collected from four datasets (e.g., sentiment analysis and code clone detection) and four LLMs (e.g., LLaMA and GPT4), we found that 1) GPT models are more confident in their outputs (e.g., 0.84 vs. 0.76 of GPT3.5 vs. CodeLLaMA), 2) LLMs are not well-calibrated and overconfident in clone detection, problem classification, and news classification tasks, and 3) existing fault detection methods perform pool on LLMs compared their performance on classical deep learning models. For example, seven out of eight methods perform worse than random selection on LLaMA. There is a need to propose methods to further enhance existing fault detection methods.
To do so, inspired by (Li et al., 2023b; Wang et al., 2021) which utilize mutation testing techniques to help test deep learning models, we propose MuCS, a prompt Mutation-based prediction Confidence Smoothing method to enhance existing fault detection methods. Specifically, we mutate the prompt by transformation methods to generate a set of prompt mutants first. Here, both text augmentation methods and code refactoring methods have been considered in MuCS. Then, we collect the output probabilities of all mutants and compute the average prediction confidence. Finally, we use the averaged confidence to perform fault detection using current fault detection methods. We compare the effectiveness of fault detection methods with and without our confidence smoothing method and found that using MuCS significantly enhances the performance of existing methods by up to 97.62%.
To summarize, the main contributions of this paper are:
This is the first study that explores the effectiveness of fault detection methods on LLMs. We found that 1) existing methods that perform well on traditional deep learning models cannot correctly detect faults in LLMs; 2) LLMs are overconfident in clone detection, problem classification, and news classification tasks; 3) the prediction confidence of LLMs is concentrated in a few intervals.
We propose MuCS, a prompt mutation-based method to smooth the confidence of LLMs and enhance the effectiveness of existing fault detection methods.
The rest of this paper is organized as follows. Section 2 reviews the related works. Section 3 presents the design of our empirical study and Section 4 summarizes the empirical results. Section 5 introduces our proposed mutation-based confidence smoothing solution and Section 6 presents its related evaluation. Section 7 discusses the limitation of this work and Section 8 concludes this work.
2. Related Work
We review the related works from the perspectives of fault detection for deep neural networks and large language model testing.
2.1. Fault Detection for Deep Neural Networks
To save the labeling budget, multiple test selection methods (Al-Qadasi et al., 2023; Zheng et al., 2023; Zhu et al., 2023a; Yin et al., 2023) have been proposed. A recent survey (Hu et al., 2024) reviewed test selection methods and divided them into fault detection methods, sampling-based model retraining methods, model selection methods, and performance estimation methods. We focus on the fault detection methods.
The early method DeepGini (Feng et al., 2020) has been proposed by Feng et al. to reveal inputs that are more likely been mispredicted by the model. DeepGini defined a new Gini score to measure the uncertainty of deep learning models on the inputs. Later on, inspired by the mutation testing in the conventional SE field, Chen et al. proposed to mutate both input samples and deep learning models and identify faults by computing the killing score. Their evaluation on more than 20 tasks demonstrated that mutation testing is useful for helping find bugs in deep learning models. Furthermore, Gao et al. proposed an adaptive test selection (ATS) method. ATS not only considers detecting faults but also tries to find diverse faults (diverse here means faults are from different categories) in deep learning models. More recently, Bao et al. first empirically proved that simple methods (i.e., output probability-based methods) perform well on fault detection. Then, they proposed a new method to compute the uncertainty scores by averaging the basic score of the input and scores from the neighbors of this input. Besides, Ma et al. conducted a comprehensive empirical study to explore the potential of test selection methods and active learning methods on fault detection. They found that MaxP is the best in terms of the correlation between the MaxP score and the misclassification. Besides the SE community, researchers from the ML community also contributed to this field. Dan et al. built the first benchmark for detecting misclassified data and found that just using the confidence score is already a strong baseline for this task. Li et al. proposed TestRank which uses graph neural networks to learn the difference between benign inputs and faults for fault detection.
Different from these works, our work is the first to explore the effectiveness of fault detection methods for large language models. We found that existing methods cannot well detect faults in LLMs and proposed the use of prompt mutation to further enhance these methods.
2.2. Large Language Model Testing
Large language models have been the most used deep learning models for most of the tasks. Comprehensively testing or evaluating the performance of LLMs becomes one of the most important directions in all the communities.
Chang et al. surveyed works that focus on evaluating LLMs (Chang et al., 2023). They classified the evaluation objectives into seven categories, natural language processing, robustness/ethics/biases/trustworthiness, social science, natural science & engineering, medical applications, agent applications, and other applications. Evaluating LLMs for code which we are more interested in lies in the category of natural science & engineering. Xu et al. (Xu et al., 2022) conducted a comprehensive study to evaluate four series of LLMs on the code generation task using the HumanEval benchmark. They found that the performance of LLMs is affected by the parameter size of models and the training time. Liu et al. (Liu et al., 2024) propose a new code synthesis evaluation framework EvalPlus to measure the correctness of code generated by LLMs. Based on this framework, they found that the existing benchmark HumanEval is not rigorous enough. The performance (passk) of LLMs drops a lot when using the new benchmark. Besides, Ma et al. (Ma et al., 2023) evaluated the capability of LLMs to understand code syntax and semantics. The experimental results show that LLMs can understand the syntax structure of code, and perform code static analysis, but cannot approximate code dynamic behavior of code.
More recently, Yuan et al. (Yuan et al., 2023) compared LLMs with conventional code models that are fine-tuned for specific code tasks. They found that in the zero-shot setting, instruction-tuned LLMs have competitive performance to fine-tuned code models. In the one-shot setting, the guidance by one-shot example sometimes harms the performance of LLMs. Du et al. (Du et al., 2024) argued that existing works that evaluate the effectiveness of LLMs for code mainly focus on simple code tasks (function-level code synthesis) and built the first class-level cod regeneration benchmark for the LLM evaluation. Based on their benchmark, they found that existing LLMs have significantly lower performance on class-level code generation tasks compared to method-level code generation tasks. There is still a big room to improve the ability of LLMs for code generation.
Unlike existing works that focus on evaluating the performance of LLMs from the perspective of preparing test cases, our work measures the capability of LLMs from another view and studies the problem of quickly identifying potential faults in LLMs to save the testing cost.
3. Empirical Study
In this work, we only focus on the problem of fault detection of LLMs. We consider both learning-based fault detection methods and output-probability-based methods in the literature. Based on the previous definition (Hu et al., 2024), given a LLM , an unlabeled test set , and a labeling budget , fault detection is to select a subset of such that where are the labels corresponding to , and is the performance measurement function. We follow the guidance of previous works (Hu et al., 2024; Ma et al., 2021; Hu et al., 2023) to select nine fault detection methods in our study. We exclude neuron coverage-based methods and surprise adequacy methods in our benchmark since the parameter size of LLMs is too big (>7B) and the coverage and adequacy score are difficult to compute.
3.1. Fault Detection Methods
Let be a set of test inputs and be the corresponding label set, and denote a given test sample and its true label, respectively. Let be the likelihood of belonging to class produced by the model .
Random Selection is the basic selection criteria that test the model using randomly selected a subset of test data.
Max Probability (MaxP) prioritizes data based on the maximum output probability . The data sample with a lower MaxP score is more likely to be a fault.
DeepGini defines a Gini score calculated by to measure the uncertainty of data. The data with higher Gini scores are treated as faults.
Entropy selects data with the minimum Shannon entropy calculated using output probabilities.
Margin measures the uncertainty of data based on the difference between the top-1 and top-2 output probabilities. A smaller difference indicates that the model is more difficult to distinguish the data between the two classes and the data should be considered faults.
Multiple-Boundary Clustering and Prioritization (MCP) contains two steps. First, it divides the data space into different areas based on the decision boundary. Here, the decision boundary is approximated by the top-2 prediction probabilities. For example, give an input and its top-2 output probabilities are and . Then this input is near the decision boundary and on the side of . MCP then selects data from different data spaces based on the score , where and are the top-1 and top-2 probabilities.
Nearest Neighbor Smoothing (NNS) first finds the neighbors and then smooths the output probabilities of each input using the outputs of neighbors. Finally, uncertainty-based fault detection methods such as DeepGini are employed to select the faults. In NNS, neighbors are searched by computing the distance between each extracted representation of the input. NNS requires the intermediate outputs of inputs to conduct data selection and, thus, cannot be used for closed-source LLMs such as GPT3.5.
Adaptive Test Selection (ATS) first projects the top-3 maximum output probabilities to the space plane. After that, it computes the coverage of each input on the plane. Then, the coverage score is used to identify if the input is a fault or not based on its difference from the coverage of the whole test set. Note that ATS can only be used for classification tasks that have greater than three categories.
TestRank initially extracts two features from the input data: 1) the output from the logits layer as intrinsic attributes, and 2) the graph information, which includes the cosine distance to other data points and the label of the data. Subsequently, a graph neural network (GNN) model is employed to assimilate the graph information and forecast the contextual attributes of the data. Finally, the contextual attributes and intrinsic attributes are amalgamated and inputted into a binary classification model to acquire proficiency in identifying failures. Comparable to the NNS method, TestRank also necessitates intermediate outputs for data selection and is not applicable to closed-source LLMs.
Except for the random selection and TestRank, all the other methods are purely based on the output probability. For the TestRank, the output probability is also a part of the information used for the detection. Thus, the prediction confidence of LLMs is important for fault detection.
3.2. Research Question
Our study plans to explore the effectiveness of existing fault detection methods on LLMs. Concretely, we answer the following two research questions:
RQ1: How confident are LLMs in their prediction? Most (seven out of nine) of the fault detection methods are purely output uncertainty-based. Therefore, checking the prediction confidence produced by LLMs before running fault detection methods is necessary.
RQ2: How effective are fault detection methods for LLMs? In this research question, we check if the existing methods demonstrated to be useful in classical deep learning models can perform well on LLMs.
3.3. Datasets and Models
Task | Class Number | Test Size | Model | Accuracy |
CodeLLaMA | 49.0% | |||
GPT3.5 | 75.5% | |||
Clone Detection (CD) | 2 | 200 | GPT4 | 77.0% |
CodeLLaMA | 17% | |||
GPT3.5 | 36% | |||
Problem Classification (PC) | 5 | 200 | GPT4 | 100.0% |
LLaMA | 89.3% | |||
GPT3.5 | 82.0% | |||
Sentiment Analysis (Sentiment) | 3 | 150 | GPT4 | 90.0% |
LLaMA | 21.5% | |||
GPT3.5 | 60.0% | |||
TagMyNews (TMN) | 7 | 200 | GPT4 | 82.0% |
In total, we consider four datasets including both natural language classification tasks and programming language classification tasks, and four types of LLMs covering both open-sourced models and closed-source models.
Clone Detection (CD): Utilizing the BigCloneBench (Svajlenko et al., 2014) dataset, we randomly selected 200 items outside of 2 classes to investigate code clone detection. Models are evaluated on their ability to predict the presence of code clones (a probability score from 0 to 1, where 0 indicates no clone and 1 indicates a clone) between code snippet pairs, a critical task for enhancing software maintenance and development practices.
Problem Classification (PC): Based on the Java250 (Puri et al., 2021) dataset from the CodeNet Project, problem classification focuses on the classification of programming challenges on LLMs. A set of 200 samples across 5 distinct problem types was chosen in a way of a one-shot learning prompt. This task tests the model’s proficiency in categorizing coding problems using a singular example for each category, requiring output as a probability distribution across classes that sum to 1.
Sentiment Analysis (Sentiment): For this task, 150 samples were extracted from the Sentiment Analysis of IMDB Movie Reviews dataset 222Autolabel, https://github.com/refuel-ai/autolabel. The challenge for models is to predict the sentiment of movie reviews, negative, neutral, and positive, ranging from 0 to 1 without any prior examples (zero-shot learning), evaluating the capacity of models for natural language understanding and emotional tone assessment.
TagMyNews (TMN): Using the TagMyNews Dataset, we randomly picked 200 articles with seven categories, which are sourced from RSS feeds of popular newspapers (Vitale et al., 2012). The task aims to assess the efficacy of the model in accurately categorizing news articles based on their textual content utilizing a one-shot learning framework. The expected output schema is similar to the problem classification task.
LLaMA and CodeLLaMA. LLaMA (Touvron et al., 2023) is known for its auto-regressive architecture aimed at chat applications. The parameters of the series of LLaMA models range from 7 billion to 70 billion and are learned from publicly available online data amounting to 2 trillion tokens. Besides, LLaMA models have been fine-tuned with a combination of supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF), focusing on aligning the models to human preferences for helpfulness and safety. CodeLLaMA (Roziere et al., 2023) is one variant of LLaMA that specifically designed for programming contexts. This adaptation involves further fine-tuning LLaMA on datasets specific to coding, thereby boosting their code-related performance capabilities. Our experiments are carried out using the meta-llama/Llama-2-7b-chat-hf and codellama/CodeLlama-7b-Instruct-hf models, available on the public Hugging Face platform.
GPT3.5 and GPT4 (Achiam et al., 2023). GPT3.5 is a fined-tuned version of GPT3 developed in January 2022. It is built on Transformer architecture while it emphasizes decoder-only mechanisms. GPT4, which is known as the state-of-the-art LLM, has a substantial leap in capability and complexity compared to GPT-3.5. Both GPT3.5 and GPT4 are closed-source LLMs, we employ gpt-3.5-turbo-0125 and gpt-4-0125-preview models in our experiments through API access, specifically within text chat application contexts.
Table 1 presents the details of our used datasets and models. We found that in some cases (marked by red), the performance of LLMs is too bad (close to random guess) and too good (with perfect accuracy). For these cases, we only evaluate their prediction confidence (RQ1) and do not conduct fault detection (RQ2) on them.
3.4. Evaluation Metrics.
In RQ1, we follow the previous work (Guo et al., 2017) to evaluate the confidence of LLMs in their prediction using metrics Prediction Confidence (Confidence) and Expected Calibration Error (ECE). Roughly speaking, Confidence is computed by probabilities associated with the predicted label. ECE first splits the range of the prediction score into equally-spaced intervals (denoted as , ) and then computes the weighted average of the bins’ accuracy/confidence. A lower ECE score indicates that the model is better calibrated, i.e., the prediction probabilities can better represent the true correctness of models.
Definition 0 (Prediction Confidence (Confidence)).
where is the number of classes.
Definition 0 (Expected Calibration Error (ECE)).
Given a test set , let be the inputs whose prediction confidence falls into the interval , be the accuracy of the inputs , and be the average confidence within , then the expected calibration error on is defined as:
In RQ2, we follow the previous work (LI et al., 2021) and evaluate the effectiveness of fault detection methods using the metric of Test Relative Coverage (TRC) which is calculated by the percentage of the number of faults divided by the labeling budget or the total number of faults whichever is minimal.
Definition 0 (Test Relative Coverage (TRC)).
Given a test set and a labeling budget , the test relative coverage on is defined as:
where represents the faults within the selected test set and represents the faults in the whole test set . A higher TRC value indicates the better performance of the fault detection method.
3.5. Configurations
Configuration of fault detection methods. NNS and TestRank need to compute the distance between data samples and require the representation of the samples. In this work, we directly use the input embeddings extracted from LLMs as the representation. Besides, TestRank is a learning-based method where training data is required for fault detection. For the clone detection, problem classification, and TagMyNews tasks, we randomly select 200 data from the original datasets as the training data. For the sentiment analysis dataset, as the original dataset only contains 200 data samples and 150 of them are used as test data, we collect the remaining 50 samples as the training data. For the data labeling process, we set the labeling budgets as 10%, 30%, and 50%. Configuration of LLMs. In our experiments, we set top_p=0.9 for LLaMA/CodeLLaMA and top_p=1 for GPT3.5/GPT4 which are the default settings used by LLMs. top_p controls the accumulation of token probabilities to select the next word for all LLMs. We set the window size of the token context as max_length= 10k for LLaMA/CodeLLaMA to match our test data’s token length. The window sizes of GPT3.5 and GPT4 are set as 16385 and 128000 respectively, using the default settings.
3.6. Implementation and Environment
For the fault detection methods, we reuse the official implementations provided by the original papers and modify them to fit our tasks. For the LLaMA and CodeLLaMA models, we use the resources provided by Hugging Face 333Hugging Face, https://huggingface.co/ in this work. For GPT3.5 and GPT4 models, we use the OpenAI’s API to access the models and collect the results. In total, it costs 267 US dollars to run GPT-related experiments. We conduct all LLaMA-related experiments on a 2.6 GHz Intel Xeon Gold 6132 CPU with an NVIDIA Tesla V100 16G SXM2 GPU.
4. Empirical Results
4.1. RQ1: Prediction Confidence
First, we check whether LLMs can confidently predict the data in our studied tasks. Figure 1 depicts the distribution (number of data in each confidence interval) of the prediction confidence produced by LLMs, where we set the number of intervals as 30. The first conclusion we can draw is that the series of GPT models are more confident with their predictions than the series of LLaMA models. On average, the confidence of LLaMA, CodeLLaMA, GPT3.5, and GPT4 on these four datasets are 0.5338, 0.6993, 0.7701, and 0.8727, respectively. Comparing the accuracy of LLMs and their average confidence (which is a notion of miscalibration (Guo et al., 2017)), we found that existing LLMs are not well-calibrated. For example, for CodeLLaMA on the clone detection task, the accuracy and average confidence are 49% and 0.758, respectively. This means CodeLLaMA is overconfident with its prediction even though it has poor performance. Overall, except for sentiment analysis, LLMs are overconfident in the other three tasks with 21.36% higher confidence than the true accuracy. Besides, ECE demonstrated that GPT4 is better calibrated than other LLMs in all our considered tasks. The Reliability diagrams (Guo et al., 2017) that visualize the model calibration can be found in Appendix A. Additionally, the results show that prediction confidence is concentrated in certain ranges for many cases. For example, for the clone detection - GPT3.5, confidence only falls into four ranges. This means the model has similar confidence to many different inputs, in turn, the output-based fault detection methods (e.g., MaxP) could produce similar uncertainty scores to these inputs.
Considering the other two cases that have poor performance reported in Table 1, CodeLLaMA for problem classification and LLaMA for TagMyNews. We found they have confused confidence on these tasks as shown in Figure 5(b) and Figure 5(c). For instance, in Figure 5(b), the most frequent confidence score is 0.5, and in Figure 5(c), there are no confidence scores that are higher than 0.8. This phenomenon indicates that CodeLLaMA and LLaMA cannot handle problem classification and news classification tasks at all.
Finally, we found that GPT4 has perfect accuracy (100%) on the problem classification task with high average confidence (0.9827 on average) and a low ECE score of 0.0143. We conjecture there is a data leakage problem of our test data, i.e., our used problem classification dataset is a part of the training data of GPT4. As there are no faults in GPT4 for the problem classification task, we exclude this model from the following fault detection study.
Answer to RQ1: LLMs are not well-calibrated and overconfident in clone detection, problem classification, and news classification tasks. Besides, prediction confidence is concentrated in a few intervals, indicating LLMs have similar confidence to most of the inputs.
4.2. RQ2: Effectiveness of Fault Detection
In this RQ, we explore the performance of existing fault detection methods on LLMs. As mentioned in Section 3.3, we exclude models that have poor performance (nearly random guesses) and perfect performance (100% accuracy) on the studied tasks.
Budget | Random | Gini | Entropy | MCP | MaxP | Margin | ATS | NNS | TestRank | |
Sentiment Analysis | 10% | 0.0667 | 0.0667 | 0.0667 | 0.1333 | 0.0000 | 0.1333 | 0.0667 | 0.0000 | 0.0667 |
30% | 0.4375 | 0.0625 | 0.1250 | 0.5000 | 0.0000 | 0.1250 | 0.1250 | 0.1250 | 0.0625 | |
50% | 0.4375 | 0.3125 | 0.2500 | 0.5625 | 0.2857 | 0.4375 | 0.3750 | 0.1875 | 0.2500 | |
Average | 0.3139 | 0.1472 | 0.1472 | 0.3986 | 0.0952 | 0.2319 | 0.1889 | 0.1042 | 0.1264 |
Budget | Random | Gini | Entropy | MCP | MaxP | Margin | ATS | ||
GPT3.5 | Clone Detection | 10% | 0.1500 | 0.3000 | 0.4500 | 0.3500 | 0.3000 | 0.3000 | - |
30% | 0.3265 | 0.4694 | 0.4490 | 0.4082 | 0.4694 | 0.4694 | - | ||
50% | 0.5102 | 0.6327 | 0.6122 | 0.6122 | 0.6327 | 0.6327 | - | ||
Problem Classification | 10% | 0.7000 | 0.8000 | 0.7500 | 0.7000 | 0.8000 | 0.8500 | 0.6500 | |
30% | 0.7167 | 0.7667 | 0.6500 | 0.6333 | 0.7833 | 0.7667 | 0.6833 | ||
50% | 0.5900 | 0.7000 | 0.6100 | 0.6400 | 0.7000 | 0.7000 | 0.6700 | ||
TagMyNew | 10% | 0.7000 | 0.5500 | 0.5500 | 0.6000 | 0.5500 | 0.5500 | 0.4000 | |
30% | 0.4000 | 0.5333 | 0.5333 | 0.4500 | 0.5500 | 0.5500 | 0.4889 | ||
50% | 0.4375 | 0.5625 | 0.5500 | 0.5000 | 0.6000 | 0.6625 | 0.8444 | ||
Sentiment Analysis | 10% | 0.0667 | 0.2667 | 0.1333 | 0.2667 | 0.2667 | 0.3333 | 0.0000 | |
30% | 0.3704 | 0.3704 | 0.4074 | 0.4444 | 0.4444 | 0.4074 | 0.2222 | ||
50% | 0.5926 | 0.6667 | 0.7037 | 0.5556 | 0.6667 | 0.7037 | 0.5185 | ||
Average | 0.4634 | 0.5515 | 0.5333 | 0.5134 | 0.5636 | 0.5771 | 0.4975 | ||
GPT4 | Clone Detection | 10% | 0.4000 | 0.5000 | 0.5000 | 0.2500 | 0.5000 | 0.2000 | - |
30% | 0.3478 | 0.5652 | 0.5435 | 0.5217 | 0.5652 | 0.3478 | - | ||
50% | 0.5435 | 0.6522 | 0.6522 | 0.6739 | 0.6522 | 0.5652 | - | ||
TagMyNews | 10% | 0.3000 | 0.4500 | 0.3500 | 0.4000 | 0.4500 | 0.4500 | 0.4000 | |
30% | 0.3333 | 0.6667 | 0.6667 | 0.5000 | 0.6389 | 0.6667 | 0.7222 | ||
50% | 0.5278 | 0.8611 | 0.8611 | 0.7778 | 0.8611 | 0.8611 | 0.8333 | ||
Sentiment Analysis | 10% | 0.0667 | 0.3333 | 0.5333 | 0.3333 | 0.3333 | 0.3333 | 0.2667 | |
30% | 0.4667 | 0.6667 | 0.7333 | 0.7333 | 0.6667 | 0.6667 | 0.3333 | ||
50% | 0.6000 | 0.9333 | 1.0000 | 0.9333 | 0.9333 | 0.9333 | 0.5333 | ||
Average | 0.3984 | 0.6254 | 0.6489 | 0.5693 | 0.6223 | 0.5582 | 0.5148 |
Table 2 and Table 5 summarize the results. The first conclusion we can draw from the results is that the existing fault detection methods cannot significantly find faults in LLMs compared to their performance on classical deep learning models (LI et al., 2021) where the best method can achieve a 0.9532 TRC score. For example, considering the detection performance on the LLaMA model, only the best method has a higher TRC score (0.3986) than random selection (0.3139). This means all the other well-designed methods perform worse than random guesses. Considering the results on GPT3.5 and GPT4 models, this phenomenon has been alleviated. All the methods perform better than random selection. However, the best methods on GPT3.5 and GPT4 have only 0.5771 and 0.6489 TRC scores. There is still a big gap between the existing performance and the idea performance, TRC=1.
Then, comparing each method, we found that no method can consistently perform better than the others. The performance of fault detection methods on LLMs highly depends on the datasets and models. However, it is interesting that when the performance of LLMs becomes better, the fault detection methods become more effective. For example, the TRC scores produced by each method on GPT4 (on average 0.6114) are much higher than the TRC scores on LLaMA (on average 0.3074). The reason is that existing methods that are based on output probabilities highly rely on the confidence of models in their outputs. As GPT4 is more confident that LLaMA, the fault detection methods can perform better on GPT4 accordingly.
Answer to RQ2: Existing fault detection methods are not effective in detecting faults in LLMs and there are no methods that can consistently perform better than the others. Methods have better performance on LLMs that have higher accuracy (GPT4) than those that have lower accuracy (GPT3.5 and LLaMA).
5. MuCS
Since the concentrated confidence will harm the performance of output-based fault detection methods (those methods will give similar uncertainty scores to the inputs), we need to diversify the prediction confidence of LLMs to enhance fault detection. It is challenging since we cannot get the internal information (e.g., output logits) of closed-source LLMs to smooth the prediction confidence. Inspired by previous works (Wang et al., 2021; Li et al., 2023b) that showed the potential of testing deep learning models by mutation analysis (one uses mutation killing score to detect faults and the other one uses mutation analysis to improve the performance of code models), we propose MuCS, a simple yet effective black-box solution to enhance fault detection by smoothing the prediction confidence from the input-level using prompt mutation.
Figure 2 presents the overall workflow of our proposed solution. It has two main steps, 1) prompt mutation, and 2) confidence smooth. Specifically, given an LLM and an input prompt p, we first utilize transformation methods to mutate the inputs and generate n mutated prompts, , , …, . For NLP tasks, we use text augmentation methods (Marivate and Sefara, 2020) (e.g., token random deletion) to transform prompts into new ones. For the code tasks, we consider both text augmentation methods and code refactoring methods (Al Dallal and Abdin, 2017) (e.g., add new Print() functions) for the prompt mutation. After that, we feed all the mutants to the LLM and collect the output confidence , ,…, produced by the LLM. Finally, a confidence smooth method is employed to compute the final confidence of the LLM on the input. In this work, we use the straightforward method, average, to smooth the confidence. Algorithm 1 summarizes our mutation-based confidence smoothing solution.
5.1. Mutation Operators
Two types of mutation operators have been used, text augmentation methods and code refactoring methods. In total, five text augmentation methods (including Synonym Replacement, Random Deletion, Random Insertion, Random Swap, and Punctuation Insertion), and four code refactoring methods (including Print Adding, Local Variable Adding, Dead If Adding, and Duplication) have been used in MuCS. Given a prompt with the length of works , each mutation operator is defined as follows.
Random Deletion (RD) randomly removes words in the with a probability. Specifically, RD assigns a random probability uniformed from 0 to 1 to each word in . Given a probability threshold , if is less than , the corresponding word will be removed. We use the default setting of MixCode (Dong et al., 2023b), in MuCS.
Random Insertion (RI) randomly selects words in first where . Then, it finds the synonyms of selected words and inserts them into the random locations of . We use the default setting of MixCode, in MuCS.
Random Swap (RW) randomly swaps two words in .
Punctuation Insertion (PI) randomly selects punctuation (e.g., !) and inserts them into the random locations of , where . We use the default setting of MixCode, in MuCS.
Print Adding (PA) randomly selects a line in the code and adds a print function with randomly generated content below it.
Local Variable Adding (LVA) randomly selects a line in the code and defines an unused local variable below it.
Dead If Adding (DIA) randomly selects a line in the code and then adds an unreachable if statement below it.
Duplication randomly copies an assignment in the code and inserts it next to this assignment.
During the mutant generation process, we randomly select mutation operators to mutate and generate its mutant .
6. Evaluation
6.1. Research Question
To check whether MuCS can help fault detection or not, we first analyze the effectiveness of existing fault detection methods using MuCS. Then, we explore the reason why MuCS enhances fault detection by analyzing its influences on the prediction confidence of LLMs.
RQ3: How effective is MuCS in enhancing fault detection? In this research question, we compare the effectiveness of fault detection methods with and without using MuCS to analyze the usefulness of our proposed solution.
RQ4: How does MuCS affect the outputs of the LLMs? We want to explore how MuCS influences the outputs of LLMs to check why it works on fault detection.
6.2. Evaluation Setup
We set the mutation times as 10 for all tasks. That is, we generate 10 mutants for each input prompt. We set (the number of mutation operators used) as 3 for all the tasks except for the clone detection. The reason is that we found the performance of LLMs on the clone detection task drops significantly (nearly random guesses, e.g., all the predicted labels of GPT3.5 are 0) when and . Therefore, we set for the clone detection task. For other tasks, LLMs have similar accuracy on the mutated datasets to the original ones (with less than 5% difference). For the other settings, we use the same settings as Section 3.
6.3. RQ3: Effectiveness of MuCS
Budget | Gini-M | Entropy-M | MCP-M | MaxP-M | Margin-M | ATS-M | NNS-M | TestRank-M | BALD |
10% | 0.0000( 100.00%) | 0.0667( 0.00%) | 0.1333( 0.00%) | 0.0000( * ) | 0.2667( 100.08%) | 0.0667( 0.00%) | 0.0000( * ) | 0.0000( 100.00%) | 0.2667 |
30% | 0.1250( 100.00%) | 0.1250( 0.00%) | 0.5000( 0.00%) | 0.2500( * ) | 0.3125( 150.00%) | 0.2500( 100.00%) | 0.0625( 50.00%) | 0.1250( 100.00%) | 0.5625 |
50% | 0.5625( 80.00%) | 0.5625( 125.00%) | 0.7500( 33.33%) | 0.5625( 96.88%) | 0.6250( 42.86%) | 0.3750( 0.00%) | 0.3750( 100.00%) | 0.5625( 125.00%) | 0.8125 |
Average | 0.2292( 26.67%) | 0.2514( 41.67%) | 0.4611( 11.11%) | 0.2708( 96.88%) | 0.4014( 97.64%) | 0.2306( 33.33%) | 0.1458( 25.00%) | 0.2292( 41.67%) | 0.5472 |
Budget | Gini-M | Entropy-M | MCP-M | MaxP-M | Margin-M | ATS-M | BALD | ||
GPT3.5 | Clone Detection | 10% | 0.4500( 50.00%) | 0.4500( 0.00%) | 0.4500( 28.57%) | 0.4500( 50.00%) | 0.4500( 50.00%) | - | 0.4500 |
30% | 0.5306( 13.04%) | 0.5102( 13.63%) | 0.5306( 29.99%) | 0.5306( 13.04%) | 0.5306( 13.04%) | - | 0.3265 | ||
50% | 0.6939( 9.67%) | 0.5510( 10.00%) | 0.5918( 3.33%) | 0.6939( 9.67%) | 0.6939( 9.67%) | - | 0.5714 | ||
Problem Classification | 10% | 0.8000( 0.00%) | 0.8000( 6.67%) | 0.5500( 21.43%) | 0.8500( 6.25%) | 0.8000( 5.88%) | 0.8500( 30.77%) | 0.8500 | |
30% | 0.8333( 8.69%) | 0.7833( 20.51%) | 0.6500( 2.64%) | 0.8500( 8.52%) | 0.8500( 10.86%) | 0.7667( 12.21%) | 0.8500 | ||
50% | 0.8000( 14.29%) | 0.8300( 36.07%) | 0.6900( 7.81%) | 0.8000( 14.29%) | 0.7700( 10.00%) | 0.7200( 7.46%) | 0.7700 | ||
TagMyNew | 10% | 0.6000( 9.09%) | 0.6500( 18.18%) | 0.4000( 33.33%) | 0.7000( 27.27%) | 0.6000( 9.09%) | 0.7000( 75.00%) | 0.7500 | |
30% | 0.6000( 12.51%) | 0.6000( 12.51%) | 0.4500( 0.00%) | 0.6167( 12.13%) | 0.5833( 6.05%) | 0.6042( 23.58%) | 0.6167 | ||
50% | 0.6875( 22.22%) | 0.6625( 20.45%) | 0.5750( 15.00%) | 0.7000( 16.67%) | 0.6750( 1.89%) | 0.7917( 6.24%) | 0.6500 | ||
Sentiment Analysis | 10% | 0.3333( 24.97%) | 0.4000( 200.08%) | 0.5333( 99.96%) | 0.2667( 0.00%) | 0.3333( 0.00%) | 0.4000( * ) | 0.3333 | |
30% | 0.5926( 59.99%) | 0.5926( 45.46%) | 0.7037( 58.35%) | 0.5556( 25.02%) | 0.6667( 63.65%) | 0.5926( 166.70%) | 0.5556 | ||
50% | 0.9259( 38.88%) | 0.8889( 26.32%) | 0.8148( 46.65%) | 0.9259( 38.88%) | 0.9259( 31.58%) | 0.6667( 28.58%) | 0.6296 | ||
Average | 0.6539( 21.95%) | 0.6432( 32.49%) | 0.5783( 19.24%) | 0.6616( 18.48%) | 0.6566( 16.66%) | 0.6769( 42.26%) | 0.6128 | ||
GPT4 | Clone Detection | 10% | 0.4500( 10.00%) | 0.4500( 10.00%) | 0.4000( 60.00%) | 0.4500( 10.00%) | 0.3000( 50.00%) | - | 0.2500 |
30% | 0.6304( 11.54%) | 0.6304( 15.99%) | 0.6304( 20.84%) | 0.6304( 11.54%) | 0.2391( 31.25%) | - | 0.2222 | ||
50% | 0.7826( 19.99%) | 0.7391( 13.32%) | 0.7826( 16.13%) | 0.7826( 19.99%) | 0.4565( 19.23%) | - | 0.4222 | ||
TagMyNews | 10% | 0.4500( 0.00%) | 0.4000( 14.29%) | 0.4000( 0.00%) | 0.4000( 11.11%) | 0.4000( 11.11%) | 0.4000( 0.00%) | 0.5000 | |
30% | 0.5556( 16.66%) | 0.5278( 20.83%) | 0.6389( 27.78%) | 0.6389( 0.00%) | 0.6389( 4.17%) | 0.7222( 0.00%) | 0.5833 | ||
50% | 0.8889( 3.23%) | 0.8056( 6.45%) | 0.7778( 0.00%) | 0.9167( 6.46%) | 0.9167( 6.46%) | 0.8611( 3.34%) | 0.7222 | ||
Sentiment Analysis | 10% | 0.4000( 20.01%) | 0.4000( 25.00%) | 0.4000( 20.01%) | 0.4000( 20.01%) | 0.4000( 20.01%) | 0.4667( 74.99%) | 0.3333 | |
30% | 0.8667( 30.00%) | 0.8667( 18.19%) | 0.8000( 9.10%) | 0.8667( 30.00%) | 0.8667( 30.00%) | 0.6667( 100.03%) | 0.5333 | ||
50% | 0.9333( 0.00%) | 0.9333( 6.67%) | 0.9333( 0.00%) | 0.9333( 0.00%) | 0.9333( 0.00%) | 0.7333( 37.50%) | 0.5333 | ||
Average | 0.6619( 6.46%) | 0.6392( 0.79%) | 0.6403( 17.09%) | 0.6687( 7.43%) | 0.5724( 4.52%) | 0.6417( 35.98%) | 0.4555 |
First, we explore whether the MuCS can enhance the existing fault detection methods or not. Before that, since the input mutation can produce multiple predictions from LLMs which is similar to the dropout uncertainty (Gal and Ghahramani, 2016). We introduce another fault detection method baseline which is based on dropout uncertainty, Bayesian Active Learning by Disagreement (BALD) (Houlsby et al., 2011).
Bayesian Active Learning by Disagreement (BALD) defines the uncertainty of an input by the disagreement of the outputs produced by LLMs on the mutants of this input. The disagreement score is calculated by , where is the predicted label of mutant and is the number of mutants. A higher BALD score indicates the input is more likely to be a fault.
Table 4 and Table 5 summarize the TRC scores of various methods enhanced by MuCS (with -M as a suffix), and the relative difference between the scores with and without using MuCS. The results demonstrate that the MuCS significantly boosts the performance of fault detection. In terms of the average TRC scores of each method on different tasks, in 19 out of 20 cases, the performance of existing fault detection methods has been increased with TRC improvements ranging from 0.0366 to 0.1794. Considering the average relative improvement, confidence smooth can enhance the performance of existing methods by up to 97.64% (LLaMA for sentiment analysis, method: Margin). Therefore, we can conclude that MuCS is effective in boosting the existing fault detection methods on LLMs.
Comparing the improvements across different LLMs, we found that improvements in GPT4 are weaker than improvements in other LLMs. On average, the improvements in LLaMA, GPT3.5, and GPT4 are 37.63%, 41.41%, and 11.78%, respectively. We conjecture that the reason for this phenomenon is that GPT4 is more powerful and more resilient to the input mutation. Thus, the smooth is less useful for GPT4. To check our assumption, we compute the average variance of the prediction output probability changes after prompt mutation of each LLM. The results show that the probability changes of GPT are negligible. For example, for the sentiment analysis task, the variance of the probability changes of LLaMA, GPT3.5, and GPT4 are 0.0047, 0.0158, and 0.0041, respectively.
Finally, even though the mutation-based confidence smooth can enhance the fault detection of LLMs, the best methods only have 0.5472, 0.6769, and 0.6687 TRC scores in LLaMA, GPT3.5, and GPT4, respectively. There is still a big room between the TRC scores achieved by existing methods and the ideal performance.
Answer to RQ3: MuCS significantly enhances the effectiveness of existing fault detection methods by up to 97.62% in terms of the TRC score.
6.4. RQ4: MuCS Affected Prediction Confidence
Model | Confidence () | ECE () | Diversity () | |||
Before | After | Before | After | Before | After | |
CD-GPT35 | 0.8535 | 0.8863 | 0.1915 | 0.2563 | 584.96 | 655.29 |
CD-GPT4 | 0.8575 | 0.8665 | 0.1630 | 0.1638 | 696.62 | 542.42 |
PC-GPT35 | 0.7659 | 0.7349 | 0.4560 | 0.3449 | 432.22 | 40.89 |
TMN-GPT35 | 0.6013 | 0.5981 | 0.0982 | 0.1030 | 205.36 | 24.69 |
TMN-GPT4 | 0.8187 | 0.7571 | 0.0407 | 0.0846 | 196.56 | 62.29 |
Sentiment-LLaMA | 0.5476 | 0.5626 | 0.3065 | 0.3041 | 92.67 | 67.00 |
Sentiment-GPT35 | 0.8596 | 0.8153 | 0.0917 | 0.1180 | 335.87 | 104.87 |
Sentiment-GP4 | 0.8320 | 0.8059 | 0.0787 | 0.1093 | 230.60 | 93.33 |
Average change | 1.78% | 4.04% | 42.67% |
We further study how MuCS affects the prediction of LLMs and whether it diversifies the prediction confidence or not. Figure 5 depicts the confidence of LLMs on the datasets before and after confidence smoothing and Table 6 summarizes the average confidence and ECE scores. The results show no big difference between the average confidence of LLMs and ECE scores before and after smoothing. This finding is consistent with the conclusion from a previous work (Zhu et al., 2023b) that states the effectiveness of fault detection methods is not directly related to the calibration of the model (most confidence calibration methods are useless or harmful for fault detection).
Then, we check whether MuCS can increase the diversity of confidence distribution or not. We compute the variance of the number of confidence scores in each range to check this diversity. The results presented in Table 6 demonstrated that, except for the clone detection tasks, LLMs have more diverse predictions on the datasets with an average 42.67% diversity improvement. Taking Figure 6(b) as an example, the confidence scores of GPT3.5 are concentrated in 1.0 which is highly contradictory to its accuracy (36%). After the mutation-based smooth, confidence scores are spread throughout the distribution which is more useful to help understand the capability of LLMs on this task.
Answer to RQ4: MuCS significantly increases the diversity of confidence distribution of LLMs with an average 42.67% improvement and eliminates the concentrated confidence problem.
7. Discussion
7.1. Limitation and Future Direction
Limitation. The potential limitations are that 1) we only consider classification tasks. Non-classification tasks such as question answering are not included in this work. The reason is that most of the existing fault detection methods are designed for classification tasks only. Even though there also exist methods that can be employed for non-classification tasks, such as PRIMA (Wang et al., 2021), applying them for LLMs is not practical. PRIMA needs to mutate deep learning models, however, as LLMs are much larger than classical deep learning models, it is difficult to mutate LLMs and save the mutants. As a result, only classification tasks are studied in our work. 2) For a given input, our proposed mutation-based confidence smooth solution needs to access LLMs times, thus, the money consumption is times than the normal way when we test closed-source LLMs.
Future direction. As LLMs (such as ChatGPT) become the most used deep learning models nowadays, the most important task people would use is the generation tasks, e.g., chatting using ChatGPT. One of the most important directions in this field is to propose fault detection methods for non-classification tasks, especially generation tasks, and then use these methods to help reveal faults in LLMs. Some metrics (Kuhn et al., 2023; Huang et al., 2023) have been designed for measuring the uncertainty in LLMs, how to combine these uncertainty metrics with existing methods could be the first try.
7.2. Threats to Validity
The internal threat lies in the implementation of fault detection methods, LLMs, and prompt mutation methods. The implementation of each fault detection method comes from the official project provided original paper. The implementation of LLMs comes from the famous open-source project Hugging Face. The code of text mutation also comes from the commonly used project TextAugment (Marivate and Sefara, 2020). The implementation of code refactoring methods modified from the previous work (Dong et al., 2023b).
The external threat comes from our studied tasks, datasets, LLMs, and fault detection methods. As mentioned in Section 7.1, we only consider the classification tasks in this work. For the tasks, we consider tasks from both NLP and SE fields. Two traditional NLP tasks and two programming tasks are included in our work. For the studied LLMs, we consider both open-source LLMs, e.g., LLaMA, and closed-source LLMs (GPT4), where GPT4 is recognized as the SOTA LLM currently. Besides, another threat that comes from the datasets and models is that there might be data leakage issues in LLMs. For example, we found GPT4 has 100% accuracy on the problem classification task which means our test set is likely in the training set of GPT4. However, it is difficult to check this issue as GPT3.5 and GPT4 are closed-source LLMs. We can only exclude the cases from the analysis of the accuracy of LLMs. For the fault detection methods, we follow the previous works (Hu et al., 2024; Ma et al., 2021; Hu et al., 2023) and explore 10 fault detection methods including both learning-based methods and output-based methods in our study. These methods are from different communities, for example, DeepGini is from the SE community and TestRank is from the ML community. We believe our studied objectives are representative and diverse enough and the conclusion drawn from this work can be generalized to other similar cases.
The construct threat could be the hyperparameter settings and the non-determinism of LLMs. We directly use the default settings suggested by Hugging Face and OpenAI’s API to build or access LLMs. Besides, we release the datasets and prompts used in our experiments to support to reproduce results reported in our work and for future research.
8. Conclusion
In this work, we first explored the potential of fault detection methods for LLMs. Based on the experiments on four tasks including NLP and code tasks, four LLMs, and nine fault detection methods, we found LLMs are overconfident with their predictions and current fault detection methods are not effective in revealing faults in LLMs. To enhance the detection performance, we proposed MuCS, which uses mutation analysis to smooth the prediction confidence of LLMs. Specifically, we utilized text transformation and code refactoring techniques to mutate the inputs and collect the averaged output probabilities of all mutants as the final prediction confidence of LLMs. Evaluation results demonstrated that MuCS significantly increased the diversity of output distribution of LLMs and enhanced the effectiveness of fault detection by up to 97.64% in terms of test relative coverage score.